# Boosting CVaR Policy Optimization with Quantile Gradients

### v1

@@ 1 Introduction
Risk avoidance is a crucial and practical consideration in sequential decision-making which inspires risk-averse reinforcement learning (RARL). Conventional risk-neutral RL Sutton et al. (1998) focuses on maximizing the total return’s expectation; in contrast, RARL involves optimizing some risk metrics of the return variable. Commonly used risk metrics include variance Tamar et al. (2012), Gini deviation Luo et al. (2023), Value-at-risk (VaR) Hau et al. (2025), conditional VaR (CVaR) Tamar et al. (2015), entropic risk measure Hau et al. (2023b), entropic VaR Su et al. (2025). In this work, we focus on (static) CVaR. Intuitively, CVaR emphasizes the worst case value of a policy’s return, i.e., the expected value under a specified quantile level $\alpha$ (also called risk level). Thus, optimizing CVaR ensures the prevention of catastrophic outcomes of a policy.
Optimizing static CVaR in RL is a non-trivial task, since it fails to decompose into per-step CVaR optimization in policy improvement Hau et al. (2023a). Therefore, dynamic programming as in risk-neutral RL cannot be applied directly. A common and practical approach is trajectory-based CVaR policy gradient (PG) using Markovian policies Tamar et al. (2015); Greenberg et al. (2022); Luo et al. (2024). CVaR-PG samples a batch of $N$ trajectories and updates the policy with $\alpha N$ of them with worst returns. This results in sample inefficiency due to 1) $1-\alpha$ portion of sampled data is discarded; 2) good-performing trajectories cannot benefit policy learning, termed blindness to success by Greenberg et al. (2022). In cases where return distribution is discrete, CVaR-PG may even suffer from vanishing gradients Greenberg et al. (2022). From its variational representation Rockafellar et al. (2000), optimizing CVaR is a bi-level problem, where the inner policy learning reduces to a risk-neutral case Bäuerle and Ott (2011). Based on this idea, several recent works aim to improve sample efficiency by assigning per-step rewards such that the expected total return of the trajectory aligns with the CVaR objective. In consequence, all trajectories are used for policy learning in a risk-neutral style, thus improving sample efficiency.
This work aims to improve sample efficiency and accelerate policy learning of CVaR-PG from another perspective. We propose augmenting it with quantile PG of tail returns. This augmentation is reasonable and intuitive, since CVaR can be regarded as the expectation of tail quantiles. The benefits of this approach are twofold. 1) Different from static CVaR, optimizing static quantile enjoys a nested form where dynamic programming is available Hau et al. (2025). Thus all data can be used for policy learning, which improves sample efficiency. 2) Good-performing trajectories are captured by the quantile PG component, which in turn overcomes the blindness to success of CVaR-PG.
@@ 2.1 Quantile, VaR and CVaR
Thus, $\mathrm{CVaR}_{\alpha}[\tilde{x}$] can be interpreted as the expected value of the bottom $\alpha$ portion of tail quantiles, i.e., $\mathrm{CVaR}_{\alpha}[\tilde{x}]=\mathbb{E}[\tilde{x}|\tilde{x}\leq\mathrm{VaR}_{\alpha}[\tilde{x}$]]. CVaR has a variational representation as Rockafellar et al. (2000)
@@ 2.2 CVaR Optimization in RL
Generally, the optimal $\alpha$-CVaR policy is history dependent Lim and Malik (2022). Consider Eq. 2 when $y$ is fixed, the policy optimization remains $\max_{\pi}-\mathbb{E}[(y-\tilde{g}_{0})^{+}]$. A common approach is to formulate an augmented MDP with new state $(s,k)\in\mathcal{S}\times\mathbb{R}$, where $k$ is a moving variable keeping track of the rewards collected so far, i.e., $k_{t}:=\sum_{i=0}^{t-1}\gamma^{i}r_{i}$ Bäuerle and Ott (2011). In this augmented MDP, the per-step reward is 0 except the terminal state that gives a reward $(y-k_{T})^{+}$. Therefore, the reward is sparse.
In practice, we may seek to learn Markovian policy given its simplicity. Prior works focusing on Markovain CVaR policy have validated its practical performance Tamar et al. (2015); Greenberg et al. (2022); Luo et al. (2024); Kim and Min (2024); Mead et al. (2025). In addition, Lim and Malik (2022) showed that if a CVaR-optimal policy is Markovian in original MDP, the optimal history-dependent policy can be recovered from the quantile value function of the Markovian policy, indicating the inter-connection between Markovian and history-dependent policies. Consider $\bar{\pi}\in\Pi_{\mathcal{M}}$ parameterized by $\theta$. Under some mild assumptions, CVaR-PG can be derived by taking derivative of Eq. 1 and is estimated by sampling $N$ trajectories $\{\tau_{i}\}_{i=1}^{N}$ using $\bar{\pi}_{\theta}$ Tamar et al. (2015), i.e.,
@@ 2.3 VaR Optimization in RL
This method differs from risk-neutral quantile-based distributional RL Dabney et al. (2018b) in that it selects the optimal action separately for each quantile level in Eq. 4, rather than optimizing the expectation over the return distribution. Alternatively, this method can be interpreted as learning the optimal value on an augmented MDP with new state $(s,\alpha)\in\mathcal{S}\times[0,1]$ and state-dependent risk-aversion.
@@ 3.1 The New VaR Bellman Operator
[…]
@@
where $\eta$ is the step size, $\partial l_{\epsilon,\alpha}^{\kappa}(\cdot):=\partial l^{\kappa}_{\max(\epsilon,\min(1-\epsilon,\alpha))}(\cdot)$ clips the $\alpha$ of $\partial l^{\kappa}_{\alpha}(\cdot)$ as defined in Eq. 6 with a small $\epsilon\in(0,\frac{1}{2})$, $\delta_{v}$ is analogous to an one-step TD error. $\tilde{u}\sim U[0,1]$ is an independent random variable. Note that when discretizing the quantile levels, e.g., as in QR-DQN Dabney et al. (2018b), $\alpha$ is guaranteed to lie in $\epsilon,1-\epsilon]$ so the clipping of $\alpha$ is never applied.
@@
Algorithm 2 Static VaR Policy Execution Hau et al. (2025)
@@ 3.2 Reducing to Markovian Policy Class
When replacing the soft quantile loss $l^{\kappa}_{\epsilon,\alpha}(\cdot)$ by the original $l_{\alpha}(\cdot)$ (or huber one), the value update is the same as quantile regression in distributional RL Dabney et al. (2018b).
@@ 3.3 Augmenting CVaR-PG with VaR-PG
Following quantile-based distributional RL Dabney et al. (2018b), we discretize the quantile levels uniformly over $[0,1]$. The deep $v$-network takes the state as input and outputs $I$ quantile values, corresponding to the quantile levels $\{\frac{1}{2}(\frac{i-1}{I}+\frac{i}{I})\}_{i=1}^{I}$. However, this architecture is known to suffer from the quantile crossing issue, where the predicted quantile values can violate the required monotonicity in the quantile level, potentially impairing policy learning Zhou et al. (2020); Luo et al. (2022). To enforce monotonicity in a simple and effective manner, we use the following mechanism. The first network output is treated as a base value, and the remaining I1I-1 outputs are interpreted as incremental deltas, i.e.,
@@
Eq. 13 is one-step advantage estimation. Since we are now working with Markovian policy, the advantage can be estimated using multi-step rollout Schulman et al. (2016) as
@@ 4 Experiments
Following Luo et al. (2024), we modify several domains such that the risk-averse policy is clear to identify to evaluate the algorithms. We start from a simple maze domain adapted from Greenberg et al. (2022), then LunarLander from the Box2D environments of OpenAI Gym Brockman et al. (2016) and InvertedPendulum from the Mujoco Todorov et al. (2012) environments of OpenAI Gym (We did not include some other MuJoCo domains considered in Luo et al. (2024), such as Ant and HalfCheetah. In these environments, agents are free to move both forward and backward, and we observe that policies may achieve high returns while wandering near the origin, which obscures meaningful notions of risk-averse behavior. The design of environments that admit well-defined and practically relevant risk-averse objectives is itself a nontrivial research problem and is beyond the scope of this paper).
Baselines. We compare our method with the original CVaR-PG (cf. Eq. 3Tamar et al. (2015), Predictive CVaR-PG (PCVaR-PG) Kim and Min (2024), Return Capping (RET-CAP) Mead et al. (2025), and the mixture policy method (denoted as MIX) in Luo et al. (2024). The method in Greenberg et al. (2022) requires the control of environment randomness, and the method in Lim and Malik (2022) is shown to not work very well in Luo et al. (2024), so these two are not included. We also include REINFORCE with baseline method Sutton et al. (1998) as a risk-neutral baseline. We denote our method as CVaR-VaR in what follows.
Settings. To ensure a relatively fair comparison, we implement all methods using on-policy PG without off-policy importance sampling (IS), thereby eliminating the potential confounding effects of IS ratio clip. The only exception is MIX, whose risk-neutral component is updated by off-policy data via offline RL (IQL Kostrikov et al. (2022)). Across all domains, methods are implemented with deep neural networks, and the policy is updated after collecting $N=20$ trajectories. Consequently, the results may differ slightly from those reported in Mead et al. (2025) and Luo et al. (2024): the former employs off-policy learning with IS ratio clip, while the latter collects a larger number of trajectories for policy updates. Additional implementation and parameter details are provided in Appendix C.
@@ A.2 VaR Bellman Optimality Equations with random reward
Note that when $\epsilon=0$, the contraction ratio will be $1$. This is the reason why we clip $\alpha\in[\epsilon,1-\epsilon]$ for $l^{\kappa}_{\alpha}$. In practice, when discretizing the quantile range, e.g. as the case in QR-DQN Dabney et al. (2018b), the quantile levels considered are guaranteed to lie in a range $\epsilon,1-\epsilon]$.
@@ C.1 Network Architecture
PCVaR-PG Kim and Min (2024) requires to learn functions taking cumulative rewards so far as input. We normalize this value to the range in $[0,1]$ and use cosine embedding as used in IQN Dabney et al. (2018a) to embed this value, i.e., denote the input of this value as $x$, denote the input size of the cosine embedding layer as $n$, denote the parameters in this linear layer as $a_{ij}$ and $b_{j}$, the $j$-th output is
@@
Our method requires to learn the quantile value function of a state. We follow QR-DQN Dabney et al. (2018b) to output $I$ values, corresponding to quantile levels $\{\frac{1}{2}(\frac{i-1}{I}+\frac{i}{I})\}_{i=1}^{I}$. To ensure the monotonicity of the quantile function, the first output value is treated as the base quantile value, and the remaining I1I-1 outputs are increasing deltas. We apply nn.functional.softplus to make delta non-negative.
@@
As mentioned, our method uses a multi-step advantage estimation. We also implement this technique for the policy learning in PCVaR-PG Kim and Min (2024) and RET-CAP Mead et al. (2025).
@@ C.2 Maze
RET-CAP Mead et al. (2025) considers $\gamma=1$ since only when $\gamma=1$, the summation of modified reward $\hat{r}_{t}=\min(k_{t},q^{*}_{\alpha})-\min(k_{t-1},q^{*}_{\alpha})$ equals $\min(R(\tau),q^{*}_{\alpha})$. So we set $\gamma=1$ and $q^{*}_{\alpha}=0$. PCVaR-PG Kim and Min (2024) requires to estimate $q^{*}_{\alpha}$, we also set this value to stabilize training, though we find the learning curves show little difference if estiamting this value during learning. The trade-off $\omega=0.5$ in CVaR-VaR.
@@ C.3 LunarLander
For RET-CAP Mead et al. (2025), $\gamma=1$. $q^{*}_{\alpha}=270\in[230,250,270,290]$. The value of $q^{*}=270$ is also suggested by the authors. For CVaR-VaR, $\omega=0.5$ for the first $40\%$ percent of iterations and linearly decay to 0 afterwards. For MIX Luo et al. (2024), the training parameters for its risk-neutral component IQL Kostrikov et al. (2022) follow the original paper.
@@ C.4 Inverted Pendulum
For RET-CAP Mead et al. (2025), $\gamma=1$. $q^{*}_{\alpha}=230\in[230,250,270,290]$. For PCVaR-PG Kim and Min (2024), $q^{*}_{\alpha}=270\in[230,250,270,290]$. For CVaR-VaR, $\omega=0.5$ for the first $25\%$ percent of iterations and is 0 afterwards. For MIX Luo et al. (2024), the training parameters for its risk-neutral component IQL Kostrikov et al. (2022) follow the original paper.

### v2

@@ 1 Introduction
Risk avoidance is a crucial and practical consideration in sequential decision-making which inspires risk-averse reinforcement learning (RARL). Conventional risk-neutral RL (Sutton et al., 1998) focuses on maximizing the total return’s expectation; in contrast, RARL involves optimizing some risk metrics of the return variable. Commonly used risk metrics include variance (Tamar et al., 2012), Gini deviation (Luo et al., 2023), Value-at-risk (VaR) (Hau et al., 2025), conditional VaR (CVaR) (Tamar et al., 2015), entropic risk measure (Hau et al., 2023b), entropic VaR (Su et al., 2025). In this work, we focus on (static) CVaR. Intuitively, CVaR emphasizes the worst case value of a policy’s return, i.e., the expected value under a specified quantile level $\alpha$ (also called risk level). Thus, optimizing CVaR ensures the prevention of catastrophic outcomes of a policy.
Optimizing static CVaR in RL is a non-trivial task, since it fails to decompose into per-step CVaR optimization in policy improvement (Hau et al., 2023a). Therefore, dynamic programming as in risk-neutral RL cannot be applied directly. A common and practical approach is trajectory-based CVaR policy gradient (PG) using Markovian policies (Tamar et al., 2015; Greenberg et al., 2022; Luo et al., 2024). CVaR-PG samples a batch of $N$ trajectories and updates the policy with $\alpha N$ of them with worst returns. This results in sample inefficiency due to 1) $1-\alpha$ portion of sampled data is discarded; 2) good-performing trajectories cannot benefit policy learning, termed blindness to success by Greenberg et al. (2022). In cases where return distribution is discrete, CVaR-PG may even suffer from vanishing gradients (Greenberg et al., 2022). From its variational representation (Rockafellar et al., 2000), optimizing CVaR is a bi-level problem, where the inner policy learning reduces to a risk-neutral case (Bäuerle and Ott, 2011). Based on this idea, several recent works aim to improve sample efficiency by assigning per-step rewards such that the expected total return of the trajectory aligns with the CVaR objective. In consequence, all trajectories are used for policy learning in a risk-neutral style, thus improving sample efficiency.
This work aims to improve sample efficiency and accelerate policy learning of CVaR-PG from another perspective. We propose augmenting it with quantile PG of tail returns. This augmentation is reasonable and intuitive, since CVaR can be regarded as the expectation of tail quantiles. The benefits of this approach are twofold. 1) Different from static CVaR, optimizing static quantile enjoys a nested form where dynamic programming is available (Hau et al., 2025). Thus all data can be used for policy learning, which improves sample efficiency. 2) Good-performing trajectories are captured by the quantile PG component, which in turn overcomes the blindness to success of CVaR-PG.
@@ 2.1 Quantile, VaR and CVaR
Thus, $\mathrm{CVaR}_{\alpha}[\tilde{x}$] can be interpreted as the expected value of the bottom $\alpha$ portion of tail quantiles, i.e., $\mathrm{CVaR}_{\alpha}[\tilde{x}]=\mathbb{E}[\tilde{x}|\tilde{x}\leq\mathrm{VaR}_{\alpha}[\tilde{x}$]]. CVaR has a variational representation as (Rockafellar et al., 2000)
@@ 2.2 CVaR Optimization in RL
Generally, the optimal $\alpha$-CVaR policy is history dependent (Lim and Malik, 2022). Consider Eq. 2 when $y$ is fixed, the policy optimization remains $\max_{\pi}-\mathbb{E}[(y-\tilde{g}_{0})^{+}]$. A common approach is to formulate an augmented MDP with new state $(s,k)\in\mathcal{S}\times\mathbb{R}$, where $k$ is a moving variable keeping track of the rewards collected so far, i.e., $k_{t}:=\sum_{i=0}^{t-1}\gamma^{i}r_{i}$ (Bäuerle and Ott, 2011). In this augmented MDP, the per-step reward is 0 except the terminal state that gives a reward $(y-k_{T})^{+}$. Therefore, the reward is sparse.
In practice, we may seek to learn Markovian policy given its simplicity. Prior works focusing on Markovain CVaR policy have validated its practical performance (Tamar et al., 2015; Greenberg et al., 2022; Luo et al., 2024; Kim and Min, 2024; Mead et al., 2025). In addition, Lim and Malik (2022) showed that if a CVaR-optimal policy is Markovian in original MDP, the optimal history-dependent policy can be recovered from the quantile value function of the Markovian policy, indicating the inter-connection between Markovian and history-dependent policies. Consider $\bar{\pi}\in\Pi_{\mathcal{M}}$ parameterized by $\theta$. Under some mild assumptions, CVaR-PG can be derived by taking derivative of Eq. 1 and is estimated by sampling $N$ trajectories $\{\tau_{i}\}_{i=1}^{N}$ using $\bar{\pi}_{\theta}$ (Tamar et al., 2015), i.e.,
@@ 2.3 VaR Optimization in RL
This method differs from risk-neutral quantile-based distributional RL (Dabney et al., 2018b) in that it selects the optimal action separately for each quantile level in Eq. 4, rather than optimizing the expectation over the return distribution. Alternatively, this method can be interpreted as learning the optimal value on an augmented MDP with new state $(s,\alpha)\in\mathcal{S}\times[0,1]$ and state-dependent risk-aversion.
@@ 3.1 The New VaR Bellman Operator
$)]\displaystyle\mathcal{T}_{\epsilon,\kappa}^{*}v(s,\alpha)=v(s,\alpha)+\eta\max_{a}\mathbb{E}[\partial l^{\kappa}_{\epsilon,\alpha}(\delta_{v}(s,\alpha,\tilde{r}(s,a),\tilde{s}^{\prime},\tilde{u}$))] (8)
@@
where $\eta$ is the step size, $\partial l_{\epsilon,\alpha}^{\kappa}(\cdot):=\partial l^{\kappa}_{\max(\epsilon,\min(1-\epsilon,\alpha))}(\cdot)$ clips the $\alpha$ of $\partial l^{\kappa}_{\alpha}(\cdot)$ as defined in Eq. 6 with a small $\epsilon\in(0,\frac{1}{2})$, $\delta_{v}$ is analogous to an one-step TD error. $\tilde{u}\sim U[0,1]$ is an independent random variable. Note that when discretizing the quantile levels, e.g., as in QR-DQN (Dabney et al., 2018b), $\alpha$ is guaranteed to lie in $\epsilon,1-\epsilon]$ so the clipping of $\alpha$ is never applied.
@@
Algorithm 2 Static VaR Policy Execution (Hau et al., 2025)
@@ 3.2 Reducing to Markovian Policy Class
When replacing the soft quantile loss $l^{\kappa}_{\epsilon,\alpha}(\cdot)$ by the original $l_{\alpha}(\cdot)$ (or huber one), the value update is the same as quantile regression in distributional RL (Dabney et al., 2018b).
@@ 3.3 Augmenting CVaR-PG with VaR-PG
Following quantile-based distributional RL (Dabney et al., 2018b), we discretize the quantile levels uniformly over $[0,1]$. The deep $v$-network takes the state as input and outputs $I$ quantile values, corresponding to the quantile levels $\{\frac{1}{2}(\frac{i-1}{I}+\frac{i}{I})\}_{i=1}^{I}$. However, this architecture is known to suffer from the quantile crossing issue, where the predicted quantile values can violate the required monotonicity in the quantile level, potentially impairing policy learning (Zhou et al., 2020; Luo et al., 2022). To enforce monotonicity in a simple and effective manner, we use the following mechanism. The first network output is treated as a base value, and the remaining I1I-1 outputs are interpreted as incremental deltas, i.e.,
@@
Eq. 13 is one-step advantage estimation. Since we are now working with Markovian policy, the advantage can be estimated using multi-step rollout (Schulman et al., 2016) as
@@ 4 Experiments
Following Luo et al. (2024), we modify several domains such that the risk-averse policy is clear to identify to evaluate the algorithms. We start from a simple maze domain adapted from Greenberg et al. (2022), then LunarLander from the Box2D environments of OpenAI Gym (Brockman et al., 2016) and InvertedPendulum from the Mujoco (Todorov et al., 2012) environments of OpenAI Gym (We did not include some other MuJoCo domains considered in Luo et al. (2024), such as Ant and HalfCheetah. In these environments, agents are free to move both forward and backward, and we observe that policies may achieve high returns while wandering near the origin, which obscures meaningful notions of risk-averse behavior. The design of environments that admit well-defined and practically relevant risk-averse objectives is itself a nontrivial research problem and is beyond the scope of this paper).
Baselines. We compare our method with the original CVaR-PG (cf. Eq. 3(Tamar et al., 2015), Predictive CVaR-PG (PCVaR-PG) (Kim and Min, 2024), Return Capping (RET-CAP) (Mead et al., 2025), and the mixture policy method (denoted as MIX) in Luo et al. (2024). The method in Greenberg et al. (2022) requires the control of environment randomness, and the method in Lim and Malik (2022) is shown to not work very well in Luo et al. (2024), so these two are not included. We also include REINFORCE with baseline method (Sutton et al., 1998) as a risk-neutral baseline. We denote our method as CVaR-VaR in what follows.
Settings. To ensure a relatively fair comparison, we implement all methods using on-policy PG without off-policy importance sampling (IS), thereby eliminating the potential confounding effects of IS ratio clip. The only exception is MIX, whose risk-neutral component is updated by off-policy data via offline RL (IQL (Kostrikov et al., 2022)). Across all domains, methods are implemented with deep neural networks, and the policy is updated after collecting $N=20$ trajectories. Consequently, the results may differ slightly from those reported in Mead et al. (2025) and Luo et al. (2024): the former employs off-policy learning with IS ratio clip, while the latter collects a larger number of trajectories for policy updates. Additional implementation and parameter details are provided in Appendix C.
@@ 5 Conclusion and Future Work
Acknowledgements
The computing resources used in this work were provided by the Digital Research Alliance of Canada (alliancecan.ca). Yudong Luo is partially supported by a FRQNT postdoctoral scholarship and IVADO. Erick Delage is partially supported by the Canadian Natural Sciences and Engineering Research Council [Grant RGPIN-2022-05261].
@@ A.2 VaR Bellman Optimality Equations with random reward
Note that when $\epsilon=0$, the contraction ratio will be $1$. This is the reason why we clip $\alpha\in[\epsilon,1-\epsilon]$ for $l^{\kappa}_{\alpha}$. In practice, when discretizing the quantile range, e.g. as the case in QR-DQN (Dabney et al., 2018b), the quantile levels considered are guaranteed to lie in a range $\epsilon,1-\epsilon]$.
@@ C.1 Network Architecture
PCVaR-PG (Kim and Min, 2024) requires to learn functions taking cumulative rewards so far as input. We normalize this value to the range in $[0,1]$ and use cosine embedding as used in IQN (Dabney et al., 2018a) to embed this value, i.e., denote the input of this value as $x$, denote the input size of the cosine embedding layer as $n$, denote the parameters in this linear layer as $a_{ij}$ and $b_{j}$, the $j$-th output is
@@
Our method requires to learn the quantile value function of a state. We follow QR-DQN (Dabney et al., 2018b) to output $I$ values, corresponding to quantile levels $\{\frac{1}{2}(\frac{i-1}{I}+\frac{i}{I})\}_{i=1}^{I}$. To ensure the monotonicity of the quantile function, the first output value is treated as the base quantile value, and the remaining I1I-1 outputs are increasing deltas. We apply nn.functional.softplus to make delta non-negative.
@@
As mentioned, our method uses a multi-step advantage estimation. We also implement this technique for the policy learning in PCVaR-PG (Kim and Min, 2024) and RET-CAP (Mead et al., 2025).
@@ C.2 Maze
RET-CAP (Mead et al., 2025) considers $\gamma=1$ since only when $\gamma=1$, the summation of modified reward $\hat{r}_{t}=\min(k_{t},q^{*}_{\alpha})-\min(k_{t-1},q^{*}_{\alpha})$ equals $\min(R(\tau),q^{*}_{\alpha})$. So we set $\gamma=1$ and $q^{*}_{\alpha}=0$. PCVaR-PG (Kim and Min, 2024) requires to estimate $q^{*}_{\alpha}$, we also set this value to stabilize training, though we find the learning curves show little difference if estiamting this value during learning. The trade-off $\omega=0.5$ in CVaR-VaR.
@@ C.3 LunarLander
For RET-CAP (Mead et al., 2025), $\gamma=1$. $q^{*}_{\alpha}=270\in[230,250,270,290]$. The value of $q^{*}=270$ is also suggested by the authors. For CVaR-VaR, $\omega=0.5$ for the first $40\%$ percent of iterations and linearly decay to 0 afterwards. For MIX (Luo et al., 2024), the training parameters for its risk-neutral component IQL (Kostrikov et al., 2022) follow the original paper.
@@ C.4 Inverted Pendulum
For RET-CAP (Mead et al., 2025), $\gamma=1$. $q^{*}_{\alpha}=230\in[230,250,270,290]$. For PCVaR-PG (Kim and Min, 2024), $q^{*}_{\alpha}=270\in[230,250,270,290]$. For CVaR-VaR, $\omega=0.5$ for the first $25\%$ percent of iterations and is 0 afterwards. For MIX (Luo et al., 2024), the training parameters for its risk-neutral component IQL (Kostrikov et al., 2022) follow the original paper.